50.00 mL of 0.125 M HCl solution was added to a 0.500 g sample of an antacid. The back titration with 0.106 M standard NaOH solution consumes 14.75 mL to reach the end-point. Calculate the effectiveness the antacid?


1
Expert's answer
2021-01-13T13:53:13-0500

The tablets of antacid contain CaCO3 that reacts with HCl according to the following equation:

CaCO3 + 2HCl "\rightarrow" CaCl2 + H2O + CO2"\uparrow" .


When an excess of HCl is added, back titration is needed to get the number of the moles of HCl reacted with antacid. This back titration is performed using NaOH:

NaOH + HCl "\rightarrow" NaCl + H2O.

As one can see, the excess moles of HCl equals the number of the moles of NaOH used to reach the end-point:

"n(HCl, excess) = n(NaOH)"

"n(HCl, excess) =cV = 0.106\cdot0.01475"

"n(HCl, excess) = 0.0015635" mol.

Then, the number of the moles that reacted with antacid is the number of the moles of HCl added to antacid minus the excess:

"n(HCl) = n(tot)-n(HCl,excess)"

"n(HCl) = cV - 0.0015635"

"n(HCl) = 0.125\cdot0.050 - 0.0015635 = 0.0046865" mol.


Using the stoichiometry of the reaction of HCl with CaCO3, we deduce that the number of the moles of CaCO3 reacted is two times less than that of HCl:

"n(CaCO_3) = \frac{n(HCl)}{2} = \frac{0.0046865}{2} = 0.002343" mol, or 4.69 mmol of active ingredient per gram of the sample.


Therefore, the mass of CaCO3 (its molar mass is 100.09 g/mol) contained in the tablet is:

"m = nM = 0.002343\cdot100.09 = 0.2345" g, or 235 mg of CaCO3 per the sample.

Therefore, the mass percent of CaCO3 in antacide sample is:

"\omega = \frac{0.2345}{0.500}\cdot100\% = 46.9\%" , or 469 mg of active ingredient per gram of the sample.